### v2

# A Matrix Model Proposal for Quantum Gravity and the Quantum Mechanics of Black Holes

Abstract
We propose a quantum mechanical theory of quantum spaces described by large $N$ noncommutative geometry as a model for quantum gravity. The theory admits fuzzy sphere as static solution. Over the fuzzy geometry, the quantum mechanics of the fermions is given by a sum of oscillators with equal frequency. The energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy. This state of the fuzzy sphere obeys the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere is identified with the black hole horizon. Moreover the set of states in the Fermi sea gives precisely the Bekenstein-Hawking entropy. We thus propose that quantum black holes are described by fuzzy spheres with a half-filled Fermi sea in our theory. We also consider a system of two fuzzy spheres by embedding them as blocks in the matrix quantum mechanics. When the distance $r$ between the two fuzzy spheres is small, the total energy of the system can be computed using perturbation theory. We show that in the leading order of large $N$ limit, the interaction energy depends on the product of the black holes masses and the Newton constant exactly as in Newton gravity. To extract the correct $r$ dependence in the long range, a resummation of all the large $N$ corrections is needed. We outline how Newton gravity for static sources may be reproduced in the long distance limit. We also show that the interaction energy is generally finite in the short distance limit, suggesting that the black hole singularity in general relativity is removed in our quantum description of black hole.
1 Introduction
Black hole arises as solution of the classical general relativity. However, a number of properties of black hole, such as the existence of an area dependent Bekenstein-Hawking entropy [1, 2], the black hole information puzzle [3, 4, 5], the existence of a black hole singularity [6], the exhibition of holography [7, 8] etc, are puzzling and it is believed that only in a theory of quantum gravity can these problems be properly understood and resolved.
A prime candidate to a theory of quantum gravity is string theory. Previously, brane models of black hole have been considered in non-perturbative string theory, with the Berkenstein-Hawking entropy reproduced successfully from a microstates counting [9, 10]. Recently progress has been achieved for black holes constructed in the AdS/CFT correspondence [11] where the Page curve behavior of the entanglement entropy of the Hawking radiation is remarkably obtained [12, 13, 14, 15]. Nevertheless, despite these success, it is highly desirable to have a direct formulation of the quantum gravity itself and be able to describe the fundamental degrees of freedom of quantized spacetime and the set of microstates explicitly without resorting to supersymmetry and duality. In this regards, large $N$ matrix model formulation of string theory, such as the BFSS matrix model [16] or the IKKT matrix model [17], is particularly appealing since in matrix model, space/time emerges from the more fundamental quantum mechanical matrix degrees of freedom in the large $N$ limit, and one can study the quantum properties of space time using the matrix quantum mechanics.
The main stream approach to quantum black hole has been a top-down one where the properties of the black hole is studied using a certain candidate theory of quantum gravity such as those mentioned above. However so far it has not been possible to study a Schwarzschild black hole this way such that the expected properties of a quantum black hole can be obtained and addressed. In a recent series of studies [18, 19, 20], we have initiated a bottom-up approach to quantum black hole by using the anticipated properties of quantum black hole as “empirical input” to guide the construction of the fundamental theory of quantum gravity. We took the assumption that quantum gravity can be formulated in terms of a quantum mechanics of fermionic and bosonic degrees of freedom, and a generic quantum mechanical model of black hole was proposed. The degrees of freedom of the quantum mechanics is supposed to take some generic form
$L=i\psi^{+}\dot{\psi}+\psi^{+}h(X)\psi-V(X),$ (1.1)
where $h(X)$ denotes some Yukawa coupling and $V(X)$ the self-interaction. Our goal was to identify basic properties of quantum gravity which are essential to the construction of the theory. Our hope was that the model would be helpful in ways similar to that of the Bohr atomic model to the building of quantum mechanics. We found [19] that if our model admits a constant density of energy eigenstates and if the Fermi sea of the system is filled up to a Fermi energy level that is inversely proportional to the system size, then the Schwarzschild radius of black hole is reproduced for the system. Moreover the system is in a highly degenerate energy state whose counting of microstates gives precisely the Bekenstein-Hawking entropy. While the result is quite encouraging, the identified conditions are sufficient ones, and it is not clear if all the assumptions can indeed be implemented consistently in a quantum mechanics. The construction of a consistent quantum mechanical theory of spacetime which includes quantum black hole is the main motivation of this work.
In this work, we propose an explicit large $N$ quantum mechanics of non-abelian bosonic and fermionic variables as a model of quantum gravity. The use of large $N$ is inspired by the previous success of the BFSS and IKKT matrix models. There are three bosonic variables $X^{a}$ in the adjoint of $SU(N)$ which correspond to the three spatial coordinates of a quantized noncommutative space. We note that the emergence of noncommutative geometry is generally expected when space (or spacetime) is quantized in quantum gravity  [^1] The studied of quantized spacetime in terms of noncommutative geometry goes back to the study of Snyder [21] and Yang [22]. . In addition, there are two fermionic variables which couple minimally to the geometry, and make up the fundamental Hilbert space of quantum gravity. We show that our theory of quantized spacetime admits fuzzy sphere as static solution. Over the fuzzy geometry, the quantum mechanics of the fermions gives a Fermi sea with uniform energy levels. It is amazing that the energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy and satisfies the the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere is identified with the black hole horizon. Moreover the set of states in the half-filled Fermi sea gives precisely the Bekenstein-Hawking entropy. We therefore propose that quantum black holes are described by fuzzy spheres with a half-filled Fermi sea in our theory. Quite amazingly, the quantum mechanics proposed here realizes pretty much all the ideas outlined earlier in [18, 19, 20].
To show that the fuzzy spheres are gravitating objects, we need to show that the interaction between them agrees with gravity, e.g. with Newton gravity in the long distance limit. We thus consider a system of two fuzzy spheres by embedding them as blocks in the matrix quantum mechanics. When the distance $r$ between the two fuzzy sphere is small, the total energy of the system can be computed using perturbation theory. We show that in the leading order of large $N$ limit, the interaction energy depends on the product of black holes masses and the Newton constant exactly as is expected in Newton gravity. We also discuss how a resummation of all the large $N$ corrections could reproduce the correct Newton limit at large distance. It is interesting to note that the interaction energy is generally finite in the short distance limit, suggesting that the black hole singularity in general relativity is removed in our quantum description of black hole.
The plan of the paper is as follows. In section 2, we give our proposal of a large $N$ quantum mechanics as a fundamental formulation of quantum gravity in 3-dimensions. In section 3, we show that the fuzzy sphere solution with a half-filled Fermi sea has the desired properties of a quantum black hole, namely with the black hole mass-radius relation and the Bekenstein-Hawking black hole entropy reproduced correctly. In section 4, we consider a system of 2 fuzzy spheres and derive their interaction energy in the small separation limit. We discuss how a large $N$ resummation of all the higher order perturbations including the off-diagonal block interaction could generate the desired gravitational forces between the quantum black holes. Discussion is found in section 5.
2 A proposed model of quantum gravity
Let us consider a theory of quantized space where the spatial coordinates becomes operators $X^{a}$ ($a=1,2,3$) and are represented by $SU(N)$ matrices. Here $N$ is large and characterizes the dimensions of the Hilbert space of quantum gravity. In addition, we propose to include fermionic degrees of freedom in the fundamental formulation of quantum gravity. In our model proposed below, they are given by the 2-components spinors $\psi^{\dagger},\psi$ in the adjoint representation of $SU(N)$. Our choice of degrees of freedom is a minimalistic one and is allowed since we do not assume supersymmetry. In fact, it is important to realize that going bottom-up, there is no physical reason to insist on having supersymmetry. We remark that in the usual top down approach, supersymmetry is often needed for the reason of consistency of formulation. e.g. for string/M theory in 10 or 11 dimensions. However, supersymmetry is irrelevant for the well-definedness and consistency of quantum mechanics as there is no issue of renormalizability here. Also the usual advantages of supersymmetric quantum mechanics such as exact computation of quantities such as index or partition function does not appeal to us since here we are after the construction of a physical theory rather than a theory that is exact or calculable. Therefore we will not insist on having supersymmetry. This allows for a much wider choice of terms in the Lagrangian which are forbidden otherwise. We propose the following definition for the quantum theory of spacetime,
$L=\mbox{tr}\left(\frac{1}{2M_{0}}\dot{X}^{a2}+\frac{M_{P}}{N^{2}}[X^{a},X^{b}]^{2}+\frac{4M_{P}}{N^{2}}X^{a2}+i\dot{\psi}^{\dagger}\psi-\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi-M_{P}{\bf 1}\right).$ (2.1)
Here $X^{a}_{mn},\psi_{mn},\psi_{mn}^{\dagger}$ are dimensionless, $\sigma^{a}$ are the Pauli matrices, $M_{0}$ and $M_{P}$ are some mass scales with $M_{P}$ being the Planck mass that is related to the Newtonian constant $G$ as
$M_{P}^{2}=\frac{2}{\pi G}.$ (2.2)
The Lagrangian includes, apart from the kinetic terms for the bosons and fermions, a Yang-Mills type potential and a Yukawa coupling. These are quite standard. In addition, there is also included a quadratic mass term. This term is forbidden by supersymmetry. In addition, we include a cosmological constant-like term. The bosonic potential of the theory is Higgs like $-\phi^{2}+\phi^{4}$. In addition to the trivial solution at the origin. There are nontrivial fuzzy spheres as nontrivial noncommutative space solutions to the theory. Due to their physical properties, we propose them as quantum description of Schwarzschild black holes.
The Lagrangian (2.1) is invariant under the global $SU(N)$ transformation
$X^{a}\to UX^{a}U^{\dagger},\quad\psi\to U\psi U^{\dagger},\quad\psi^{\dagger}\to U\psi^{\dagger}U^{\dagger},\qquad UU^{\dagger}=1.$ (2.3)
The time derivative should be promoted to a covariant derivative if the $SU(N)$ symmetry becomes a time-dependent one. The theory also has an $SO(3)$ rotational invariance where $X^{i}$ transforms as a vector and $\psi$ in the spinor representation. The quantum mechanical Hamiltonian is given by
$H=\mbox{tr}\left(\frac{M_{0}}{2}P^{i2}-\frac{M_{P}}{N^{2}}[X^{a},X^{b}]^{2}-\frac{4M_{P}}{N^{2}}X^{a2}+\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi+M_{P}{\bf 1}\right).$ (2.4)
where $P^{a}=\dot{X}^{a}/M_{0}$ is the conjugate momentum for the bosonic degrees of freedom and satisfies the commutation relation
$[X^{a}_{mk},P^{b}_{nl}]=i\delta^{ab}\delta_{mn}\delta_{kl}.$ (2.5)
As for the fermionic degrees of freedom, the conjugate momentum $\pi=i\psi^{\dagger}$ give rises to the anticommutation relation
$\{\psi^{\dagger}_{mk\alpha},\psi_{nl\beta}\}=\delta_{mn}\delta_{kl}\delta_{\alpha\beta},\quad\alpha,\beta=1,2.$ (2.6)
We note that in the large $N$ limit, our model is equivalent to some non-supersymmetric version of the IKKT-like instantonic model in 4 dimensions by a large $N$ reduction [23].
We propose that the large $N$ quantum mechanics (2.4) provides a fundamental formulation of quantum gravity in 3 space dimensions. We remark that our proposal of a large $N$ matrix quantum mechanics as a fundamental theory of quantum gravity is similar in philosophy to that of the M(atrix) theory [24]. However we do not assume supersymmetry, and so we can write down our model directly for 3 space dimensions, and we can include a mass term. We note that a mass term has been considered in the IKKT matrix model where interesting classical solutions such as fuzzy spacetime [25] and expanding universe [26, 27, 28, 29], have been obtained. However, our proposal of a large $N$ quantum mechanics as a fundamental formulation of quantum gravity and the connection with black hole are new.
3 Quantum black hole as fuzzy sphere with half-filled Fermi sea
The classical equation of motion for a bosonic matrix configuration is given by
$-\frac{1}{M_{0}}\ddot{X}^{a}+\frac{4M_{P}}{N^{2}}\left([X^{b},[X^{a},X^{b}]]+2X^{a}\right)=0.$ (3.1)
For static configuration, this becomes
$[[X^{a},X^{b}],X^{b}]=2X^{a}.$ (3.2)
This can be solved by the spin $j=(N-1)/2$ representation of $SU(2)$, which are given by $2j+1=N$ dimensional matrices satisfying
$[X^{a},X^{b}]=i\epsilon^{abc}X^{c}.$ (3.3)
Due to the Casimir relation
$\sum_{a}X^{a2}=\frac{N^{2}-1}{4}{\bf 1}$ (3.4)
the configuration (3.3), (3.4) define a fuzzy sphere. Let us introduce dimensional coordinates $Y^{a}=2l_{P}X^{a}$ for some length scale $l_{P}$, the solution becomes
$[Y^{a},Y^{b}]=\frac{2iR}{\sqrt{N^{2}-1}}\epsilon_{abc}Y^{c},\quad\sum_{a}Y^{a2}=R^{2}{\bf 1},$ (3.5)
where $R^{2}=(N^{2}-1)l_{P}^{2}$. This describes a fuzzy sphere of radius $R$. In the following, we will take $l_{P}$ to be the Planck length given by
$l_{P}=\frac{2}{\pi M_{P}}=\sqrt{\frac{2G}{\pi}}.$ (3.6)
We are interested in the energy of the fuzzy sphere system. Over the geometry of the fuzzy sphere, the Hamiltonian of the system is given by
$H=H_{B}+H_{F},$ (3.7)
where
$H_{B}=\frac{NM_{P}}{2}(1+\frac{1}{N^{2}})$ (3.8)
obtained from the constant, $X^{2}$ and $X^{4}$ terms in (2.4), is the energy arises from the fuzzy sphere geometry, and
$H_{F}=\frac{M_{P}}{N^{2}}\psi^{\dagger}\sigma^{a}X^{a}\psi$ (3.9)
is the fermionic part of the Hamiltonian. It is easy to see that $H_{F}$ describes a collection of fermionic oscillators with equal frequency. To show this, we note that for the fuzzy sphere, the matrix $K:=\sigma^{a}X^{a}$ satisfies $K^{2}-K-\frac{N^{2}-1}{4}{\bf 1}=0$ and so it has eigenvalues $(N+1)/2$ or $-(N-1)/2$. Since it is traceless, therefore $N-1$ of the eigenvalues are positive and $N+1$ of them are the negative ones. In the leading order of large $N$ limit, $K$ has the eigendecomposition
$K_{(m\alpha)(n\beta)}=\frac{N}{2}\;\sum_{p=1}^{N}\left({\cal U}^{p}_{m\alpha}{\cal U}^{p\dagger}_{n\beta}-{\cal V}^{p}_{m\alpha}{\cal V}^{p\dagger}_{n\beta}\right).$ (3.10)
Here ${\cal U}^{p}_{n\beta},{\cal V}^{p}_{n\beta},p=1,\cdots,N$ are eigenvectors of $K$ with positive and negative eigenvalue. Let us from now on consider the leading large $N$ limit. Introduce the fermionic oscillators
$\xi^{p}_{k}:={\cal U}^{p\dagger}_{n\beta}\psi_{nk\beta},\quad\chi^{p\dagger}_{k}:={\cal V}^{p\dagger}_{n\beta}\psi_{nk\beta},$ (3.11)
which satisfy
$\{\xi^{p}_{k},\xi^{q\dagger}_{l}\}=\delta^{pq}\delta_{kl},\quad\{\chi^{p}_{k},\chi^{q\dagger}_{l}\}=\delta^{pq}\delta_{kl},$ (3.12)
then $H_{F}$ takes on the form as claimed above:
$H_{F}=\frac{M_{P}}{2N}\left(\sum_{p,k=1}^{N}\xi_{k}^{p\dagger}\xi_{k}^{p}+\chi_{k}^{p\dagger}\chi_{k}^{p}-N^{2}\right).$ (3.13)
$H_{F}$ has the eigenstates
$\left|\Psi^{p_{1}\cdots p_{r}q_{1}\cdots q_{s}}_{k_{1}\cdots k_{r}l_{1}\cdots l_{s}}\right\rangle:=\xi^{p_{1}\dagger}_{k_{1}}\cdots\xi^{p_{r}\dagger}_{k_{r}}\chi^{q_{1}\dagger}_{l_{1}}\cdots\chi^{q_{s}\dagger}_{l_{s}}|0\rangle,$ (3.14)
where $|0\rangle$ is the fermionic Fock vacuum defined by
$\xi^{p}_{k}|0\rangle=\chi^{q}_{l}|0\rangle=0,\quad\forall p,q,k,l$ (3.15)
and the eigenvalues
$H_{F}=\frac{M_{P}}{2N}n,\quad n:=r+s-N^{2},$ (3.16)
where $n=-N^{2},\cdots,0,\cdots,N^{2}$ specifies the energy level within the Fermi sea. The lowest level of energy is $-NM_{P}/2$ which corresponds to an empty Fermi sea, while the highest energy $NM_{P}/2$ corresponds to an completely filled Fermi sea. In the leading order of large $N$ limit, the fuzzy sphere has the radius
$R=Nl_{P}$ (3.17)
and the total Hamiltonian $H$ can be written as
$H=\frac{M_{P}}{2N}\sum_{a=1}^{N^{2}}\xi_{a}^{\dagger}\xi_{a}+\chi_{a}^{\dagger}\chi_{a},$ (3.18)
where we have labeled the oscillators with $a:=(p,k)$. Physically, we may give an interpretation to (3.18) by saying that the fuzzy sphere has been divided into $N^{2}$ cells with unit cell area of $\Delta A=4\pi l_{P}^{2}$ of the Planck size. Each cell, whose location is labeled by $a=(p,k)$, is populated by a pair of oscillators $\xi_{a},\chi_{a}$ which describes the quantum fluctuations over the fuzzy sphere. This is precisely the description of “partons” suggested in Susskind’s holographic picture [8] of the black hole. It is quite amusing that this conjectured feature of holography get a concrete realization in our quantum mechanics of spacetime.
Let us consider the $n=0$ energy state which corresponds to having half of the oscillator levels in the Fermi sea filled. In this case, the total energy (3.18) of the fuzzy sphere system is
$E=\frac{NM_{P}}{2}.$ (3.19)
Using the relations (3.6) and (3.17), this becomes
$E=\frac{R}{2G}.$ (3.20)
This is precisely the Schwarzschild mass-radius relation if the fuzzy sphere is identified with the black hole horizon and the total energy of the system is identified with the mass $M$ of the black hole
$E=M.$ (3.21)
We note that the relation (3.21) holds if we assume the principle of equivalence of internal (non-gravitational) energy and gravitational mass. In the next section, we will consider multiple fuzzy spheres configuration in our model and show that the interaction energy depends of the product $GM_{1}M_{2}$ precisely as in Newton’s gravity if the gravitational mass given by (3.21). This gives a simple explanation of the equivalence principle from the quantum mechanics.
Next let us consider the microstates counting. In fact, the level $n$ eigenvalue has a degeneracy of
$\Omega_{n}=\binom{2N^{2}}{N^{2}+n},$ (3.22)
which corresponds to the ways to fill exactly $N^{2}+n$ of the oscillator levels. For the $n=0$ energy state, we have
$\Omega_{0}=2^{2N^{2}}$ (3.23)
in the leading order of large $N$. These microstates of the system at the energy (3.19) give rises to the entropy $S=\log_{2}\Omega_{0}$:
$S=2N^{2}=\frac{A}{4G},$ (3.24)
which is precisely the Bekenstein-Hawking entropy of a Schwarzschild black hole. It is remarkable that our quantum mechanical theory (2.4) admits solution that matches precisely the desired properties of a quantum black hole. We thus propose that a quantum black hole is described in our theory by a fuzzy sphere geometry in the energy state $n=0$. We note that the state $n=0$ is singled out by the fact that $\Omega_{n}$ is maximized at $n=0$. It is amazing that a quantum black hole in our description is characterized to be one, given the mass of the black hole is fixed, having the maximal allowed amount of microstates. In the next section, we provide further justification by showing that the fuzzy spheres interact with each other in consistent with Newton gravity in the large distance.
Finally let us comment on the Bekenstein entropy bound [30] which set the maximal amount of entropy that can be contained within a region of space with radius $R$ and energy $E$:
$S\leq 2\pi RE.$ (3.25)
Although counter examples of this bound is known, it has been proven in quantum field theory [31]. Presumably it also holds in a consistent theory of quantum gravity. Let us check it against our theory. Indeed the bound is saturated for the $n=0$ state of the fuzzy sphere. For positive $n$, the inequality is satisfied since $S(n)<S(0)=2\pi RE_{0}<2\pi RE_{n}$. However, it is easy to see that the bound is violated for negative $n$. Physically, it means the fuzzy spheres in $n=0$ state (black hole) and the positive $n$ more energetic state (presumably represent stars and compact objects) are allowed in the theory, while those states with negative $n$ should be excluded. It is possible that these states becomes unstable in the higher order perturbation theory. It would be interesting to understand this better.
4 Multiple black holes and Newton gravity
The reproduction of the static properties of a quantum black hole is a strong evidence that the fuzzy sphere solution does describe a quantum black hole and that our theory (2.4) is a creditable proposal. Below we provide further justification to our proposal by showing that the fuzzy spheres interact in manner that is characteristic of Newton gravity.
Let us consider block diagonal configuration of the following form
$X^{a}=\left(\begin{array}[]{c|c}X_{1}^{a}&0\\ \hline\cr 0&X^{a}_{2}\end{array}\right),$ (4.1)
where $X_{\rm i}^{a}$ is of dimensions $N_{\rm i}\times N_{\rm i}$, ${\rm i}=1,2$ and $N=N_{1}+N_{2}$. We consider large $N_{\rm i}$ such that the ratios $\alpha_{\rm i}:=N_{\rm i}/N$ are fixed. The matrices $X^{a}$ are traceless, but $X^{a}_{\rm i}$ are generally not. Let us introduce the coordinates
$x^{a}_{\rm i}:=\frac{1}{N_{\rm i}}\mbox{tr}X_{\rm i}^{a},$ (4.2)
which satisfies
$N_{1}x_{1}^{a}+N_{2}x_{2}^{a}=0,$ (4.3)
due to the tracelessness of $X^{a}$. We note that since the mass of the isolated black hole is given by
$M_{\rm i}=\frac{N_{\rm i}M_{P}}{2},$ (4.4)
the relation (4.3) allows $x^{a}_{\rm i}$ to be interpreted as the location of the black holes with respect to the center of mass of the system. As a result, $X^{a}_{\rm i}$ can be decomposed into it’s trace and the traceless part $X_{\rm i}^{0a}$ as
$X^{a}_{\rm i}=x_{\rm i}^{a}+X_{\rm i}^{0a}$ (4.5)
and the Hamiltonian of the theory is given by
$H=H_{1}+H_{2}+H_{12},$ (4.6)
plus kinetic terms for $X^{0a}_{\rm i}$ and $x_{\rm i}^{a}$. Here
$H_{\rm i}:=H_{\rm i}^{0}+H_{\rm i}^{F},\quad{\rm i}=1,2$ (4.7)
where the term
$H_{\rm i}^{0}:=\alpha_{\rm i}^{2}\mbox{tr}_{\rm i}\left(-\frac{M_{P}}{N_{\rm i}^{2}}[X_{\rm i}^{0a},X_{\rm i}^{0b}]^{2}-\frac{4M_{P}}{N_{\rm i}^{2}}X_{\rm i}^{0a2}+\frac{M_{P}}{N_{\rm i}^{2}}\mbox{tr}_{\rm i}\psi^{\dagger}\sigma^{a}X_{\rm i}^{0a}\psi\right)+N_{\rm i}M_{P}$ (4.8)
is obtained from the non-abelian part $X_{\rm i}^{0a}$, while the fermionic term
$H^{F}_{\rm i}:=\frac{M_{P}}{N^{2}}\sum_{m,n=1}^{N_{\rm i}}\psi^{\dagger}_{mn}\sigma^{a}x_{\rm i}^{a}\psi_{mn}$ (4.9)
and the bosonic term
$H_{12}:=-\frac{4M_{P}}{N^{2}}(N_{1}x_{1}^{2}+N_{2}x_{2}^{2})$ (4.10)
are obtained from the abelian $U(1)$ part $x^{a}_{\rm i}$ of the configuration. In the above, the trace $\mbox{tr}_{\rm i}$ is taken over the $N_{\rm i}\times N_{\rm i}$ subspace. Finally the terms $H_{\rm i}^{F},H_{12}$ represent interaction between the two blocks and depend on their separation distance (in dimensionless matrix units)
$\Delta x:=\sqrt{(x_{1}^{a}-x_{2}^{a})^{2}}.$ (4.11)
This comes because (4.3) implies that $x_{1}^{a}=\alpha_{2}(x_{1}^{a}-x_{2}^{a})$ etc. As a result, the fermionic interaction term $H^{F}_{\rm i}$ is proportional to $\Delta x$, and the bosonic interaction term is proportional to $\Delta x^{2}$
$H_{12}=-\frac{4M_{P}}{N}\alpha_{1}\alpha_{2}\Delta x^{2}.$ (4.12)
In general, solutions to the equation of motion of (4.6) are not static. Nevertheless, one can consider static configuration and use it to derive the static potential between the blocks. More general velocity dependent forces can also be derived by considering time dependent configuration. Therefore let us consider the equation of motion in the static limit. Note that $H^{0}_{\rm i}$ is identical to the Hamiltonian for a single black hole except for a replacement of $M_{P}\to\alpha_{\rm i}^{2}M_{P}$. As a result, we obtain the same equation of motion and $X^{0a}_{\rm i}$ is solved by fuzzy sphere with radius
$R_{\rm i}=N_{\rm i}l_{P}.$ (4.13)
The Hamiltonian $H_{\rm i}^{0}$ admits the eigenstates
$\left|\Psi^{p_{1}\cdots p_{r}q_{1}\cdots q_{s}}_{k_{1}\cdots k_{r}q_{1}\cdots q_{s}}\right\rangle:=\xi^{p_{1}\dagger}_{k_{1}}\cdots\xi^{p_{r}\dagger}_{k_{r}}\chi^{q_{1}\dagger}_{l_{1}}\cdots\chi^{q_{s}\dagger}_{l_{s}}|0\rangle,$ (4.14)
where $\xi^{p}_{k},\chi^{p}_{l}$ are defined as in (3.11) using the eigendecomposition (3.10) for $\sigma^{a}X_{\rm i}^{0a}$ and the indices $m,n$ range over the block of $X_{\rm i}$. Consider the set of states where half of the oscillators over the fuzzy sphere are excited (i.e. $r+s=N_{\rm i}^{2}$), the fuzzy sphere has the energy
$H^{0}_{\rm i}=\frac{M_{P}N_{\rm i}}{2}(2-\alpha_{\rm i}^{2}).$ (4.15)
The coarse graining of this ensemble of microstates gives the Bekenstein-Hawking entropy $S_{\rm i}=2N_{\rm i}^{2}=A_{\rm i}/4G$.
Next let us consider the interaction term $H^{a}_{\rm i}$. For small $\Delta x/N$, we can consider $H^{a}_{\rm i}$ to be a perturbation and use the perturbation theory for degeneracy eigenstates to determine its correction to the energy. This requires the knowledge of the matrix elements $\langle i|H^{F}_{\rm i}|j\rangle$ and it’s eigenvalues $\lambda_{p}$. Here $|i\rangle,i=1,\cdots,\Omega_{0}:=2^{2N_{\rm i}^{2}}$ are the unperturbed states (4.14) with $r+s=N_{\rm i}^{2}$. This is however not only impossible to do as it involves too large a number of states, but it is also unnecessary. A more meaningful quantity is the ensemble average of the corrections,
$\frac{1}{\Omega_{0}}\sum_{p}\lambda_{p}$ (4.16)
We show in the appendix that this vanishes. As a result, the total energy of the two fuzzy spheres system is given by
$E=\frac{NM_{P}}{2}(1+3\alpha_{1}\alpha_{2})-\frac{4M_{P}}{N}\alpha_{1}\alpha_{2}\Delta x^{2}.$ (4.17)
Here we have used $\alpha_{1}^{3}+\alpha_{2}^{3}=1-3\alpha_{1}\alpha_{2}$. The second term is a correction to the first term in the leading ordering of small $\Delta x^{2}/N^{2}$. In principle, one should include the higher order perturbation as well as the interaction term from the off diagonal blocks of (4.1). By summing over all these contributions, one can expect the final result for the system energy to be a function of $\Delta x/N$ whose leading term in the small $\Delta x/N$ expansion gives (4.17). In other words, the large-$N$ summation would result in the energy for the two fuzzy spheres system
$E=\frac{NM_{P}}{2}(1+5\alpha_{1}\alpha_{2})-NM_{P}\alpha_{1}\alpha_{2}f(\frac{\Delta x}{N}),$ (4.18)
where $f$ is some function with the property that $f(x)=1+4x^{2}+O(x^{3})$ for small $x$. Defining the coordinate distance between the two black holes as $r=\Delta x\,l_{P}$, we have $H=M_{1}+M_{2}+5\mu-\frac{GM_{1}M_{2}}{r}g(r),$ where $M_{\rm i}=N_{\rm i}M_{P}/2$ is the mass of the black hole in isolation, $\mu=M_{1}M_{2}/(M_{1}+M_{2})$ is the reduced mass, and $g(r):=\frac{4r}{R}f(\frac{r}{R})$. Note that the factor $GM_{1}M_{2}$ appears exactly the way one expects for Newton gravity. One can define the interaction energy as $V(r):=H(r)-H(r=\infty)=-\frac{GM_{1}M_{2}}{r}g(r)$. Note that it is generally true that
$V(0)=-\frac{4GM_{1}M_{2}}{R}.$ (4.19)
This is independent of the form of $f$. The removal of the infinite interaction energy at the origin is generic and is a prediction of our proposed theory of quantum spacetime. Note also that the Newton gravity
$V(r)=-\frac{GM_{1}M_{2}}{r}$ (4.20)
will be recovered in the leading order of large distance $r/R=\Delta x/N\gg 1$ if $f$ is such that $g=1+O(1/r)$. An example is $f(x)=(1-x)/(1-x-4x^{2})$. In this case, $g(r)=(r/R-1)/(r/R+\frac{1}{4}(R/r-1))$. It is important to devise method to reliably compute the potential $V(r)$ between the fuzzy spheres in order to confirm that gravity does emerge from our proposed theory of quantized spacetime. More generally it is important to understand how a description involving metric and Einstein gravity emerges in the low energy approximation.
5 Discussion
In this paper we have proposed a theory of quantum space and gravity as a large $N$ quantum mechanics of non-abelian bosonic and fermionic coordinates. The quantum mechanics has static solution whose bosonic part is given by a fuzzy sphere. Over the fuzzy sphere geometry, the fermionic part of the theory is given by a collection of fermionic oscillators all with the same frequency. The energy state where exactly half of the Fermi sea is filled contains the maximal amount of degeneracy. This half-filled fuzzy sphere observes the mass-radius relation of a Schwarzschild black hole if the fuzzy sphere radius is identified with the horizon size. Moreover, the coarse graining of the set of quantum states in the Fermi sea gives precisely the Bekenstein-Hawking entropy. As a result, we propose that a quantum black hole is described by a half-filled fuzzy sphere in our theory.
We have also considered interaction between these fuzzy spheres by including the quantum mechanical perturbation arisen from a separation of them in the matrix space. We find that the interaction energy between the fuzzy spheres has a dependence on the masses and Newton constant exactly as in the Newton gravity. We conjecture that when the full large $N$ resummation of all the higher order perturbation effects including the off-diagonal block terms is performed, the exact result will capture the correct gravitational interaction between the static quantum black holes, with the Newton gravity reproduced in the large distance limit. The obtained modified gravity may be relevant for the dark matter and dark energy problem.
In this paper we have considered static solutions. It would be interesting to consider non-static solution in order to capture the relativistic effects. Our proposed quantum mechanics is for quantum space. Rotational black hole should be included in our model. It is interesting to understand whether and how matter would arise in our quantum mechanics. In particular, how gauge interaction can be incorporated and how a charged back hole can be obtained.
We note that the change of coordinates $x^{\mu}\to x^{\prime}{}^{\mu}(x)$ in general relativity is replaced in our description by the unitary transformation
$X^{i}\to UX^{i}U^{-1}.$ (5.1)
It is important to understand how the diffeomorphism symmetry of general relativity emerges from the quantum mechanics in some classical limit, and how Einstein gravity emerges and get modified in our theory. Einstein has believed that geodesic equation should not be taken as an independent assumption but derived from the field equations for empty space. This program can be implemented using our quantum mechanics. For example, one may consider the motion of a light probe block in the presence of a heavy one ($N_{1}\ll N_{2}$). In some limit, we may ignore the back reaction of the light block on the heavy one and take the heavy block as a background. It is possible that one may encode the effect of the heavy block on the probe in terms of some effective metric and effective energy-momentum tensor, and the motion of the probe in the background metric in terms of some effective geodesic motion.
In the general analysis of AMPS [32] concerning firewalls and complementarity, it was concluded that the following three statements cannot all be true: (i) Hawking radiation is in a pure state, (ii) the information carried by the radiation is emitted from the region near the horizon, with low energy effective field theory valid beyond some microscopic distance from the horizon, and (iii) the infalling observer encounters nothing unusual at the horizon. Our study of the two fuzzy sphere systems show that nothing singular occurs as the distance between them decreases to zero. Taking one of them as a probe, it means no fire wall and no black hole singularity. On the other hand, since the black hole horizon is described by a noncommutative fuzzy geometry in our model, the existence of a low energy effective theory is questionable due to UV/IR mixing [33] (see also UV/IR mixing in fuzzy sphere [34]) and modified causality and non-locality of noncommutative field theory [35]. It would be interesting to study these points further. In our description, the fuzzy sphere is dynamical in general and it will be interesting to develop some kind of effective membrane description for the horizon, and to compare it with the classical membrane paradigm [36].
Acknowledgments
The support of this work by NCTS and the grant 110-2112-M-007-015-MY3 of the National Science and Technology Council of Taiwan is gratefully acknowledged.
Appendix A Ensemble average of 1st order perturbation
In this appendix, we use perturbation theory to compute the average correction to the energy of the system due to the interaction term between the center coordinates of the fuzzy sphere with the fermions.
In general, consider adjoint fermions $\psi_{mn}$ of $SU(N)$ and the Hamiltonian
$h_{0}=\psi^{\dagger}\sigma^{a}X^{0a}\psi,$ (A.1)
where $X^{0a}$ is given by a fuzzy sphere. In the large $N$ limit, $h_{0}$ can be written as a sum of fermionic oscillators with the same frequency $\omega$,
$h_{0}=\frac{N}{2}\left(\sum_{p,k=1}^{N}(\xi_{k}^{p\dagger}\xi_{k}^{p}+\chi_{k}^{p\dagger}\chi_{k}^{p})-N^{2}\right),$ (A.2)
where the fermionic oscillators
$\xi_{k}^{p}:={\cal U}^{p\dagger}_{n\beta}\psi_{nk\beta},\quad\chi_{k}^{p\dagger}:={\cal V}^{p\dagger}_{n\beta}\psi_{nk\beta}$ (A.3)
are defined using the eigenvectors ${\cal U}^{p},{\cal V}^{p}$ of the matrix $K:=\sigma^{a}X^{0a}$:
$(\sigma^{a}X^{0a})_{(m\alpha)(n\beta)}=\frac{N}{2}\sum_{p=1}^{N}\left({\cal U}^{p}_{m\alpha}{\cal U}^{p\dagger}_{n\beta}-{\cal V}^{p}_{m\alpha}{\cal V}^{p\dagger}_{n\beta}\right).$ (A.4)
They satisfy the orthonormality condition
${\cal U}^{p\dagger}_{m\alpha}{\cal U}^{q}_{m\alpha}={\cal V}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\alpha}=\delta^{pq},\qquad{\cal U}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\alpha}=0$ (A.5)
and the completeness relation
${\cal U}^{p}_{m\alpha}{\cal U}^{p\dagger}_{n\beta}+{\cal V}^{p}_{m\alpha}{\cal V}^{p\dagger}_{n\beta}=\delta_{mn}\delta_{\alpha\beta}.$ (A.6)
As a result, $h_{0}$ has the eigenstates
$\left|\Psi^{p_{1}\cdots p_{r}q_{1}\cdots q_{s}}_{k_{1}\cdots k_{r}l_{1}\cdots l_{s}}\right\rangle:=\xi^{p_{1}\dagger}_{k_{1}}\cdots\xi^{p_{r}\dagger}_{k_{r}}\chi^{q_{1}\dagger}_{l_{1}}\cdots\chi^{q_{s}\dagger}_{l_{s}}|0\rangle,$ (A.7)
where $|0\rangle$ is the Fock vacuum and the eigenvalues
$h_{0}=n,$ (A.8)
where $n:=r+s-N^{2}=-N^{2},\cdots,0,\cdots N^{2}$ is the energy level within the Fermi sea.
We are interested in the correction to the energy state of $h_{0}$ due to term of the form
$h^{\prime}:=\sum_{m,n=1}^{N}\psi^{\dagger}_{mn}\sigma^{a}x^{a}\psi_{mn}.$ (A.9)
We will treat $h^{\prime}$ as a perturbation. The unperturbed level $n$ has a degeneracy of $\Omega_{n}=\binom{2N^{2}}{n+N^{2}}$. Denoting the set of unperturbed eigenstates by
$\mathbb{V}_{n}:=\{|i\rangle,i=1,\cdots,\Omega_{n}\},$ (A.10)
where $|i\rangle$ are those states (A.7) with $r+s=n+N^{2}$. With the perturbation $h^{\prime}$ turns on, some of these states will be lifted. Rather then trying to determine the splitting $\lambda_{p}$ for each individual states, it is more meaningful to compute the average correction to the level $n$ eigenvalue,
$\overline{h^{\prime}}:=\frac{1}{\Omega_{n}}\sum_{p=1}^{\Omega_{n}}\lambda_{p}=\frac{1}{\Omega_{n}}\sum_{|i\rangle\in\mathbb{V}_{n}}\langle i|h^{\prime}|i\rangle.$ (A.11)
We claim that $\overline{h^{\prime}}=0$. To see this, let us use
$\psi_{mn\alpha}={\cal U}^{p}_{m\alpha}\xi^{p}_{n}+{\cal V}^{p}_{m\alpha}\chi^{p\dagger}_{n}$ (A.12)
to rewrite $h^{\prime}$ as
$h^{\prime}=P^{pq}\xi^{p\dagger}_{k}\xi^{q}_{k}+Q^{pq}\chi^{p}_{k}\chi^{q\dagger}_{k}+T^{pq}\xi^{p\dagger}_{k}\chi^{q\dagger}_{k}+{\rm h.c.}$ (A.13)
where the matrices $P^{pq}$ etc are given by
$P^{pq}:={\cal U}^{p\dagger}_{m\alpha}{\cal U}^{q}_{m\beta}\sigma^{a}_{\alpha\beta}x^{a},\quad Q^{pq}:={\cal V}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\beta}\sigma^{a}_{\alpha\beta}x^{a},\quad$ $T^{pq}:={\cal U}^{p\dagger}_{m\alpha}{\cal V}^{q}_{m\beta}\sigma^{a}_{\alpha\beta}x^{a}.$ (A.14)
It is easy to see that the $T$ terms do not contribute to the matrix element and so
$\langle i|h^{\prime}|i\rangle=N\binom{N^{2}-1}{r}\binom{N^{2}}{s}\times\mbox{tr}P-N\binom{N^{2}-1}{s}\binom{N^{2}}{r}\times\mbox{tr}Q$ (A.15)
for any state $|i\rangle$ of the form (A.7) with $r$ and $s$ fixed. Using $\sum_{r}\binom{N^{2}-1}{r}\binom{N^{2}}{N^{2}-r}=\Omega_{0}/2$,the sum over all states of form (A.7) with a fixed level $n=0$ is given by
$\sum_{|i\rangle\in\mathbb{V}_{0}}\langle i|h^{\prime}|i\rangle=\frac{N}{2}\Omega_{0}\mbox{tr}(P-Q).$ (A.16)
Using the definitions (A.14) and (A.4), we obtain
$\overline{h^{\prime}}=\mbox{tr}(K\sigma^{a}X^{0a}).$ (A.17)
This vanishes since $\mbox{tr}(X^{0a})=0$ and hence our claim that $\overline{h^{\prime}}=0$.